# Ideals, Gröbner Bases, and PCPs

### v1

@@ 5 Zero-on-Variety Test
In this subsection, we discuss an efficient test to decide whether a given oracle vanishes on a subset of points/variety. Let $V\subset\mathbb{F}_{q}^{m}$ be a set with a Gröbner generating set $\mathfrak{G}(V)$ with extension degree $d_{V}$ and complexity $k$ (see Section 4 for formal definitions). Let $P:\mathbb{F}_{q}[x_{1},\ldots,x_{m}]\to\mathbb{F}_{q}$ be a polynomial of degree $d$ and say $d\leq d_{V}$. Informally, the main goal of this section is:
@@
1 Input: Degree parameter $d$, subset $V\subseteq\mathbb{F}_{q}^{m}$, Gröbner generating set $\mathfrak{G}(V)$,
string $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m+k}$, $\bm{\alpha}\in\mathbb{F}_{q}^{m}$, element $t\in\mathbb{F}_{q}^{\times}$,
and oracle access to $(f,\mathcal{M},\mathcal{M}^{\prime})$ where $f\in(\mathbb{F}_{q})^{q^{m}}$, $\mathcal{M}\in(\mathbb{F}_{q})^{q^{m+k}}$, and $\mathcal{M}^{\prime}\in(\mathbb{F}_{q}^{d})^{q^{2(m+k)}}$.
@@
There exists an absolute constant $C>0$ such that for every $d,q\in\mathbb{N}$ satisfying $q>Cd^{3}$, for every subset $V\subset\mathbb{F}_{q}^{m}$ with extension degree $\leq d$ and Gröbner complexity $k$, the following holds. The standard verifier $\mathcal{ZERO}$ satisfies the following properties:
@@
$\Rightarrow\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{LC}^{\mathcal{M},\mathcal{M}^{\prime}}_{d}((\bm{\alpha},\mathbf{0});\mathbf{a},t)\neq\mathcal{R}(\bm{\alpha},\mathbf{0})]\;\leq\;0.08+\dfrac{d}{q-1}.$ (4)
@@
$\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{R}(\bm{\alpha},\mathbf{0})\neq 0]\;\leq\;0.12+\dfrac{d}{q-1}$
$\iff\quad\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{R}(\bm{\alpha},\mathbf{0})=0]\;\geq\;0.88-\dfrac{d}{q-1}.$
@@
$\Pr_{\bm{\alpha}}[\mathcal{R}(\bm{\alpha},\mathbf{0})=0]\;\geq\;0.88-\dfrac{d}{q-1}.$
We choose $C$ in the statement of Lemma 5.3 large enough such that $0.88-\frac{d}{q-1}>\frac{d}{q}$. The polynomial distance lemma (Theorem 3.1) then implies that $\mathcal{R}(\mathbf{x},\mathbf{0})\equiv 0$. This finishes the proof of 5.4. ∎
@@
$\Rightarrow\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{LC}^{\mathcal{M},\mathcal{M}^{\prime}}_{d}((\bm{\alpha},\varphi(\bm{\alpha}));\mathbf{a},t)=\mathcal{R}(\bm{\alpha},\varphi(\bm{\alpha}))]\;\geq\;0.92-\dfrac{d}{q-1}$
$\Rightarrow\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{LC}^{\mathcal{M},\mathcal{M}^{\prime}}_{d}((\bm{\alpha},\varphi(\bm{\alpha}));\mathbf{a},t)\neq R(\bm{\alpha})]\;\leq\;0.08+\dfrac{d}{q-1}.$ (6)
Recall that $P\neq R$ and from the polynomial distance lemma (Theorem 3.1), we have:
@@
$\leq\;\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{LC}^{\mathcal{M},\mathcal{M}^{\prime}}_{d}((\bm{\alpha},\varphi(\bm{\alpha}));\mathbf{a},t)\neq R(\bm{\alpha})]\,+\,\Pr_{\bm{\alpha}}[R(\bm{\alpha})=f[\bm{\alpha}]]\quad\leq\quad 0.08+\dfrac{2d^{2}}{q-1}+\delta.$
@@
$\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{LC}^{\mathcal{M},\mathcal{M}^{\prime}}_{d^{2}}((\bm{\alpha},\varphi(\bm{\alpha}));\mathbf{a},t)\neq f[\bm{\alpha}]]\;\geq\;0.91-\delta\;\geq\;0.04,$
@@
From Lemma 4.7 and 4.8, we know that $\mathcal{V}_{\mathsf{PCP}}$ can compute the Gröbner generating set of complexity $k$ in time $q^{\mathcal{O}(m)}$. Let $E:V\times V\to\{0,1\}\subset\mathbb{F}_{q}$ be the edge function for the input graph $G=(V,E)$, defined as $E(u,v)=1$ if and only if $(u,v)\in E$. Let $\widehat{E}:\mathbb{F}_{q}^{m}\times\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$ denote an extension of $E$ and from the second item of Lemma 4.3, we know that $\deg(\widehat{E})\leq 2d$. By Lemma 4.6 and 4.8, the standard verifier $\mathcal{V}_{\mathsf{PCP}}$ can compute both, the extension degree $d$ and the extension $\widehat{E}$ in time $q^{\mathcal{O}(m)}$.
For simplicity, we will describe a standard verifier $\mathcal{V}$ for $3$-$\mathsf{COLOR}$ and then $\mathcal{V}_{\mathsf{PCP}}$ will be repeating $\mathcal{V}$ for $\mathcal{O}(1)$ times. More particularly, the standard verifier $\mathcal{V}$ will have soundness $\gamma$ for some absolute constant $\gamma\in(0,1)$, i.e. $\mathcal{V}$ rejects with probability at least $\gamma$. The standard verifier $\mathcal{V}_{\mathsf{PCP}}$ will repeat $\mathcal{V}$ for $\mathcal{O}(1/\gamma)$ times and return Reject if any one of the iterations return Reject. As $\mathcal{O}(1/\gamma)=\mathcal{O}(1)$, the number of random bits, queries, and running time of $\mathcal{V}_{\mathsf{PCP}}$ are a constant factor multiple of the number of random bits, queries, and running time of $\mathcal{V}$ respectively. So for rest of the proof, it will be sufficient to describe a standard verifier $\mathcal{V}$ which uses $c^{\prime}(k+m)\log q$ random bits, makes $\ell^{\prime}$ queries to proofs over alphabets of size $c^{\prime}\cdot(d\log q)$, have soundness guarantee of $\gamma$, and has running time $q^{\mathcal{O}(m+k)}$, for some constants $c^{\prime},\ell^{\prime},$ and $\gamma\in(0,1)$. From the previous paragraph, we know that $\mathcal{V}_{\mathsf{PCP}}$ can compute the Gröbner generating set, extension degree $d$, and extension $\widehat{E}$, all in time $q^{\mathcal{O}(m)}$. So we will assume that our standard verifier $\mathcal{V}$ has access to all of them.
@@
    Let [^5] Recall $k$ is the Gröbner complexity of $\mathfrak{G}(V)$. $\mathcal{M}_{A}:\mathbb{F}_{q}^{m+k}\to\mathbb{F}_{q}$ denote a vanishing certificate polynomial for $A|_{V}$ (see Definition 5.1 for a formal definition). Let $d_{1}:=\deg(\mathcal{M}_{A})\leq\deg(A)\leq 3d$. Let $\mathcal{M}_{A,\mathrm{lines}}^{(3d)}$ be the $(3d)^{th}$ lines table for $\mathcal{M}_{A}$.
@@
1 2 Input: Degree parameter $d$, subset $V$, Gröbner generating set for $V$, polynomial $\widehat{E}$,
strings $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m}$, $\bm{\alpha},\bm{\beta}\in\mathbb{F}_{q}^{2m}$, $\bm{\gamma}_{1},\bm{\gamma}_{2}\in\mathbb{F}_{q}^{m+k}$, $\bm{\mu}_{1},\bm{\mu}_{2}\in\mathbb{F}_{q}^{2(m+k)},$ $t\in\mathbb{F}_{q}^{\times}$,
@@
Since both the polynomials in the above inequality have degree $\leq(3d)$ and because $3d/q<9\gamma$, the polynomial distance lemma (Theorem 3.1) implies Equation 9. An analogous argument shows Equation 10. This finishes the proof of 6.4. ∎
@@
Recall that $\widetilde{\chi}$ is $\varepsilon$-far from any degree $d$ extension of a proper $3$-coloring of $G$. In particular, for $\gamma<\varepsilon/4$, we know that $P_{\widetilde{\chi}}$ is not an extension of a proper $3$-coloring. In particular, there exists vertices $\mathbf{u},\mathbf{v}$ such that $(\mathbf{u},\mathbf{v})\in E$ and $P_{\widetilde{\chi}}(\mathbf{u})=P_{\widetilde{\chi}}(\mathbf{v})$. In other words, $P_{\widetilde{B}}$ does not vanish on $V\times V$. From Lemma 5.3, we know that event $\mathcal{E}_{4}$ happens with probability $\geq(0.04-4\gamma)>\gamma$. Hence we have shown that either one of $\mathcal{E}_{1},\mathcal{E}_{2}$, or $\mathcal{E}_{3}$ happens with probability $\gamma$, otherwise $\mathcal{E}_{4}$ happens with probability $\gamma$. This finishes the proof of 6.3.
@@ 7 The PCP Theorem with One Composition
More precisely, we need an extension of our PCP to a Robust Assignment-Tester (or equivalently a Robust PCP of Proximity [BGHSV]). The definition below is due to Dinur and Reingold [DinurReingold]. We will assume throughout this section that there is a single growing parameter $n$ and all other parameters ($R,q,\varepsilon,\ldots$) are (possibly constant) functions of $n.$
@@
Recall that two functions $f,g:S\rightarrow T$ are said to be $\delta$-close if they differ on at most a $\delta$-fraction of their inputs and $\delta$-far if they differ on at least a $\delta$-fraction of their inputs.
Definition 7.1 (Robust Assignment-Testers, combining Definitions 3.1 and 3.4 from [DinurReingold]).
A Robust Assignment-Tester with parameters $(R,s,\ell,\delta,\varepsilon,w,\rho)$ is a reduction whose input is a Boolean circuit $\varphi$ of size $n$ over Boolean variables $X$. The reduction runs in time $\operatorname{poly}(n,R)$ and outputs a system of $R(n)$ Boolean circuits $\psi=\{\psi_{1},...,\psi_{R}\}$, each $\psi_{i}$ of size at most $s(n)$ over Boolean variables $X$ and auxiliary variables $Y$ such that the following conditions hold.
  • Each $\psi_{i}$ depends on a set of $\ell$ variables from $X\cup Y$, which we denote $\mathrm{Vars}(\psi_{i})$. The variables in Y take values in an alphabet $\Sigma$, and are accessible to $\psi_{i}$ as a tuple of $w=\lceil\log_{2}(|\Sigma|)\rceil$ bits.
@@
      Completeness: If $\sigma$ satisfies the input circuit $\varphi$, then there exists an assignment $\tau:Y\rightarrow\Sigma$ such that $\sigma\cup\tau$ satisfies all of $\psi_{1},\ldots,\psi_{R}.$
@@
      Robust Soundness: If $\sigma$ is $\delta$-far from any satisfying assignment of the input circuit $\varphi$, then for any assignment $\tau:Y\rightarrow\Sigma$ and for any $(1-\varepsilon)$-fraction of the $\psi_{i}$’s, the restricted assignment $(\sigma\cup\tau)|_{\mathrm{Vars}(\psi_{i})}$ is $\rho$-far from any satisfying assignment of $\psi_{i}$.
A (not necessarily robust) Assignment-Tester with parameters $(R,s,\ell,\delta,\varepsilon,w)$ is defined as above, except that we replace the Robust Soundness condition with the following weaker Soundness condition.
  • Soundness: If $\sigma$ is $\delta$-far from any satisfying assignment of $\varphi$, then for any assignment $\tau:Y\rightarrow\Sigma$ and for a $(1-\varepsilon)$ fraction of the $\psi_{i}$’s, the restricted assignment $(\sigma\cup\tau)|_{\psi_{i}}$ is not a satisfying assignment of $\psi_{i}$.
[DinurReingold] showed that the existence of one assignment-tester with specific parameters is enough to prove the PCP theorem. We state the observation formally below.
Observation 7.2 (Section 3.1 in [DinurReingold]).
@@
it suffices to show that there is an assignment-tester with parameters $R(n)=\operatorname{poly}(n)$, $s,\ell,w\in\mathcal{O}(1)$ and $\delta,(1-\varepsilon)\in\Omega(1).$
@@
We will prove the existence of such an assignment-tester by composing two robust assignment-testers with suitable parameters. The process of composition starts with two robust assignment-testers $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ and produces a robust assignment-tester $\mathcal{A}_{3}$ with many parameters (e.g. $s,\ell,w$) that are dictated by the parameters of $\mathcal{A}_{2}$. The robust assignment-tester $\mathcal{A}_{3}$ tests assignments to the Boolean circuit $\varphi$ that is an input to $\mathcal{A}_{1}$.
@@
The construction is simple: The tester $\mathcal{A}_{3}$ first runs $\mathcal{A}_{1}$ with input $\varphi$ and then runs $\mathcal{A}_{2}$ on each of the circuits produced by $\mathcal{A}_{1}.$ The high-level idea behind it is that we test an assignment to the input circuit $\varphi$ to $\mathcal{A}_{1}$ by testing that each of the ‘local views’ of the various circuits produced by $\mathcal{A}_{1}$ is close to a satisfying assignment, where the latter process is carried out by the circuits produced by reduction $\mathcal{A}_{2}.$ This leads us to the following observation of Dinur and Reingold [DinurReingold, Lemma 3.5], which we need only as a black box.
Lemma 7.3 (Composing robust assignment-testers: Lemma 3.5 in [DinurReingold]).

### v2

@@ 5 Zero-on-Variety Test
In this subsection, we discuss an efficient test to decide whether a given oracle vanishes on a subset of points/variety. Let $V\subset\mathbb{F}_{q}^{m}$ be a set with a Macaulay basis $\mathfrak{G}(V)$ with extension degree $d_{V}$ and complexity $k$ (see Section 4 for formal definitions). Let $P:\mathbb{F}_{q}[x_{1},\ldots,x_{m}]\to\mathbb{F}_{q}$ be a polynomial of degree $d$ and say $d\leq d_{V}$. Informally, the main goal of this section is:
@@
1 Input: Degree parameter $d$, subset $V\subseteq\mathbb{F}_{q}^{m}$, Macaulay basis $\mathfrak{G}(V)$,
randomness $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m+k}$, $\bm{\alpha}\in\mathbb{F}_{q}^{m}$, element $t\in\mathbb{F}_{q}^{\times}$, and
oracle access to $(f,\mathcal{M},\mathcal{M}^{\prime})$ where $f\in(\mathbb{F}_{q})^{q^{m}}$, $\mathcal{M}\in(\mathbb{F}_{q})^{q^{m+k}}$, and $\mathcal{M}^{\prime}\in(\mathbb{F}_{q}^{d})^{q^{2(m+k)}}$.
@@ 1 Input: Degree parameter $d$, subset $V\subseteq\mathbb{F}_{q}^{m}$, Macaulay basis $\mathfrak{G}(V)$,
There exists an absolute constant $C>0$ such that for every $d,q\in\mathbb{N}$ satisfying $q>Cd^{3}$, for every subset $V\subset\mathbb{F}_{q}^{m}$ with extension degree $\leq d$ and Macaulay complexity $k$, the following holds. The standard verifier $\mathcal{ZERO}$ satisfies the following properties:
@@
$\Rightarrow\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{LC}^{\mathcal{M},\mathcal{M}^{\prime}}_{d}((\bm{\alpha},\mathbf{0});\mathbf{a},t)\neq\mathcal{R}(\bm{\alpha},\mathbf{0})]\;\leq\;0.8+\dfrac{d}{q-1}.$ (4)
@@
$\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{R}(\bm{\alpha},\mathbf{0})\neq 0]\;\leq\;0.84+\dfrac{d}{q-1}$
$\iff\quad\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{R}(\bm{\alpha},\mathbf{0})=0]\;\geq\;0.16-\dfrac{d}{q-1}.$
@@
$\Pr_{\bm{\alpha}}[\mathcal{R}(\bm{\alpha},\mathbf{0})=0]\;\geq\;0.16-\dfrac{d}{q-1}.$
We choose $C$ in the statement of Lemma 5.3 large enough such that $0.16-\frac{d}{q-1}>\frac{d}{q}$. The polynomial distance lemma (3.1) then implies that $\mathcal{R}(\mathbf{x},\mathbf{0})\equiv 0$. This finishes the proof of 5.4. ∎
@@
$\Rightarrow\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{LC}^{\mathcal{M},\mathcal{M}^{\prime}}_{d}((\bm{\alpha},\varphi(\bm{\alpha}));\mathbf{a},t)=\mathcal{R}(\bm{\alpha},\varphi(\bm{\alpha}))]\;\geq\;0.2-\dfrac{d}{q-1}$
$\Rightarrow\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{LC}^{\mathcal{M},\mathcal{M}^{\prime}}_{d}((\bm{\alpha},\varphi(\bm{\alpha}));\mathbf{a},t)\neq R(\bm{\alpha})]\;\leq\;0.8+\dfrac{d}{q-1}.$ (6)
Recall that $P\neq R$ and from the polynomial distance lemma (3.1), we have:
@@
$\leq\;\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{LC}^{\mathcal{M},\mathcal{M}^{\prime}}_{d}((\bm{\alpha},\varphi(\bm{\alpha}));\mathbf{a},t)\neq R(\bm{\alpha})]\,+\,\Pr_{\bm{\alpha}}[R(\bm{\alpha})=f[\bm{\alpha}]]\quad\leq\quad 0.8+\dfrac{2d^{2}}{q-1}+\delta.$
@@
$\Pr_{\bm{\alpha},\mathbf{a},t}[\mathcal{LC}^{\mathcal{M},\mathcal{M}^{\prime}}_{d^{2}}((\bm{\alpha},\varphi(\bm{\alpha}));\mathbf{a},t)\neq f[\bm{\alpha}]]\;\geq\;0.19-\delta\;\geq\;0.04,$
@@
From Lemma 4.9 and 4.10, we know that $\mathcal{V}_{\mathsf{PCP}}$ can compute the Macaulay basis of complexity $k$ in time $q^{\mathcal{O}(m)}$. Let $E:V\times V\to\{0,1\}\subset\mathbb{F}_{q}$ be the edge function for the input graph $G=(V,E)$, defined as $E(u,v)=1$ if and only if $(u,v)\in E$. Let $\widehat{E}:\mathbb{F}_{q}^{m}\times\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$ denote an extension of $E$ and from the second item of Lemma 4.4, we know that $\deg(\widehat{E})\leq 2d$. By Lemma 4.8 and 4.10, the standard verifier $\mathcal{V}_{\mathsf{PCP}}$ can compute both, the extension degree $d$ and the extension $\widehat{E}$ in time $q^{\mathcal{O}(m)}$.
For simplicity, we will describe a standard verifier $\mathcal{V}$ for $3$-$\mathsf{COLOR}$ and then $\mathcal{V}_{\mathsf{PCP}}$ will be repeating $\mathcal{V}$ for $\mathcal{O}(1)$ times. More particularly, the standard verifier $\mathcal{V}$ will have soundness $\gamma$ for some absolute constant $\gamma\in(0,1)$, i.e. $\mathcal{V}$ rejects with probability at least $\gamma$. The standard verifier $\mathcal{V}_{\mathsf{PCP}}$ will repeat $\mathcal{V}$ for $\mathcal{O}(1/\gamma)$ times and return Reject if any one of the iterations return Reject. As $\mathcal{O}(1/\gamma)=\mathcal{O}(1)$, the number of random bits, queries, and running time of $\mathcal{V}_{\mathsf{PCP}}$ are a constant factor multiple of the number of random bits, queries, and running time of $\mathcal{V}$ respectively. So for rest of the proof, it will be sufficient to describe a standard verifier $\mathcal{V}$ which uses $c^{\prime}(k+m)\log q$ random bits, makes $\ell^{\prime}$ queries to proofs over alphabets of size $c^{\prime}\cdot(d\log q)$, have soundness guarantee of $\gamma$, and has running time $q^{\mathcal{O}(m+k)}$, for some constants $c^{\prime},\ell^{\prime},$ and $\gamma\in(0,1)$. From the previous paragraph, we know that $\mathcal{V}_{\mathsf{PCP}}$ can compute the Macaulay basis, extension degree $d$, and extension $\widehat{E}$, all in time $q^{\mathcal{O}(m)}$. So we will assume that our standard verifier $\mathcal{V}$ has access to all of them.
@@
    Let [^5] Recall $k$ is the Macaulay complexity of $\mathfrak{G}(V)$. $\mathcal{M}_{A}:\mathbb{F}_{q}^{m+k}\to\mathbb{F}_{q}$ denote a vanishing certificate polynomial for $A|_{V}$ (see Definition 5.1 for a formal definition). Let $d_{1}:=\deg(\mathcal{M}_{A})\leq\deg(A)\leq 3d$. Let $\mathcal{M}_{A,\mathrm{lines}}^{(3d)}$ be the $(3d)^{th}$ lines table for $\mathcal{M}_{A}$.
@@
1 2 Input: Degree parameter $d$, subset $V$, Macaulay basis for $V$, polynomial $\widehat{E}$,
@@
Since both the polynomials in the above inequality have degree $\leq(3d)$ and because $3d/q<9\gamma$, the polynomial distance lemma (3.1) implies Equation 9. An analogous argument shows Equation 10. This finishes the proof of 6.4. ∎
@@
Recall that $\widetilde{\chi}$ is $\varepsilon$-far from any degree $d$ extension of a proper $3$-coloring of $G$. In particular, for $\gamma<\varepsilon/4$, we know that $P_{\widetilde{\chi}}$ is not an extension of a proper $3$-coloring. In particular, there exists vertices $\mathbf{u},\mathbf{v}$ such that $(\mathbf{u},\mathbf{v})\in E$ and $P_{\widetilde{\chi}}(\mathbf{u})=P_{\widetilde{\chi}}(\mathbf{v})$. In other words, $P_{\widetilde{B}}$ does not vanish on $V\times V$. From Lemma 5.3, we know that event $\mathcal{E}_{4}$ happens with probability $\geq 0.04>\gamma$. Hence we have shown that either one of $\mathcal{E}_{1},\mathcal{E}_{2}$, or $\mathcal{E}_{3}$ happens with probability $\gamma$, otherwise $\mathcal{E}_{4}$ happens with probability $\gamma$. This finishes the proof of 6.3.
@@ 7 The PCP Theorem with One Composition
More precisely, we need an extension of our PCP to a Robust Assignment-Tester (or equivalently a Robust PCP of Proximity [5]). The definition below is due to Dinur and Reingold [9] with minor changes. In particular, our notion of robust soundness is defined for every distance parameter instead of fixed quantities. We will assume throughout this section that there is a single growing parameter $n$ and all other parameters ($r,\ell,a$) are (possibly constant) functions of $n.$
@@
Recall that two functions $f,g:S\rightarrow T$ are said to be $\delta$-close if they differ on at most a $\delta$-fraction of their inputs and $\delta$-far if they differ on at least a $\delta$-fraction of their inputs. For a predicate $A$ and an input $x$ in its domain, we define $\delta(x,A^{-1}(1))$ to be $\min_{y:A(y)=1}\delta(x,y)$.
Definition 7.1 (Robust Assignment-Testers, combining Definitions 3.1 and 3.4 from [9]).
A Robust Assignment-Tester with parameters $(r,\ell,a,\rho)$ is a reduction whose input is a Boolean circuit $\varphi$ of size $n$ over Boolean variables $X$. Letting $R=R(n):=2^{r(n)}$, the reduction runs in time $\operatorname{poly}(n,R)$ and outputs a system of $R$ Boolean circuits $\psi=\{\psi_{1},...,\psi_{R}\}$, each $\psi_{i}$ of size at most $a(n)$ over Boolean variables $X$ and auxiliary variables $Y$ such that the following conditions hold.
  • Each $\psi_{i}$ depends on a set of $\ell$ variables from $X\cup Y$, which we denote $\mathrm{Vars}(\psi_{i})$. The variables in Y take values in an alphabet $\Sigma$, and are accessible to $\psi_{i}$ as a tuple of $\lceil\log_{2}(|\Sigma|)\rceil$ bits [^6] In particular, this means that $\lceil\log_{2}(|\Sigma|)\rceil\leq a(n)$..
@@
      Completeness: If $\sigma$ satisfies the input circuit $\varphi$, then there exists an assignment $\tau:Y\rightarrow\Sigma$ such that $\sigma\cup\tau$ satisfies all of $\psi_{1},\ldots,\psi_{R}$, i.e.: Pr_i[R][ψ_i((στ)—_Vars(ψ_i))=1]=1.
@@
      Robust Soundness: For every assignment $\tau:Y\rightarrow\Sigma$, it holds that E_i[R][δ((στ)—_Vars(ψ_i), ψ_i^-1(1))] ρδ(σ, φ^-1(1)). [^7] Here the relative Hamming distance $\delta()$ is defined over bits since $\psi_{i}$ is defined to be a Boolean circuit.
A (not necessarily robust) Assignment-Tester with parameters $(r,\ell,a,\rho)$ is defined as above, except that we replace the Robust Soundness condition with the following weaker Soundness condition.
  • Soundness: For every assignment $\tau:Y\rightarrow\Sigma$, it holds that
    $\Pr_{i\in[R]}[\psi_{i}((\sigma\cup\tau)|_{\mathrm{Vars}(\psi_{i})})\neq 1]\geq\rho\cdot\delta(\sigma,\varphi^{-1}(1)).$
[9] showed that the existence of one assignment-tester with specific parameters is enough to prove the PCP theorem. We state the observation formally below, adjusting to our definition.
Observation 7.2 (Section 3.1 in [9]).
@@
it suffices to show that there is an assignment-tester with parameters $r(n)=\mathcal{O}(\log n)$, $\ell=\mathcal{O}(1),a=\mathcal{O}(1)$ and $\rho=\Omega(1).$
@@
We will prove the existence of such an assignment-tester by composing two robust assignment-testers with suitable parameters. The process of composition starts with two robust assignment-testers $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ and produces a robust assignment-tester $\mathcal{A}_{3}$ with parameters that are dictated by the parameters of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. The robust assignment-tester $\mathcal{A}_{3}$ tests assignments to the Boolean circuit $\varphi$ that is an input to $\mathcal{A}_{1}$.
@@
The construction is simple: The tester $\mathcal{A}_{3}$ first runs $\mathcal{A}_{1}$ with input $\varphi$ and then runs $\mathcal{A}_{2}$ on each of the circuits produced by $\mathcal{A}_{1}.$ The high-level idea behind it is that we test an assignment to the input circuit $\varphi$ to $\mathcal{A}_{1}$ by testing that each of the ‘local views’ of the various circuits produced by $\mathcal{A}_{1}$ is close to a satisfying assignment, where the latter process is carried out by the circuits produced by reduction $\mathcal{A}_{2}.$ This leads us to the following observation of Dinur and Reingold [9, Lemma 3.5].
Lemma 7.3 (Composing robust assignment-testers: Lemma 3.5 in [9]).